[Brahms-l] Shift report 20050427 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Apr 27 2005 - 16:09:44 EDT
Shifter leader was kh
Shifters were 


Beam until 13:00 when the dumped on schedule.  After that there were several problems with the Cyrogenic system which prevented more beam.
Runs achieved:
Run  tr2   tr3
14305 1539  130k
14306 3395  564k
14307 2054  166k
14308 1643   92k

This brings the total tr2 for FS 8 deg 2442 to 27k and
for tr3 to 9.38M if we take out the runs from 14221 through 14250 
where CC1-5 were not working.

Changed D5 polarity to A

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 27 16:09:58 2005

This archive was generated by hypermail 2.1.8 : Wed Apr 27 2005 - 16:10:18 EDT